OpenStudy (anonymous):

Evaluate the following: 4^2 ÷ (3 + 1). My answer is 2, am i right?

jimthompson5910 (jim_thompson5910):

No that is not correct

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes I'm sure

OpenStudy (anonymous):

i keep getting the same answer when i solve i..

jimthompson5910 (jim_thompson5910):

oh it's not 42, you changed it to 4^2

jimthompson5910 (jim_thompson5910):

I see what you did there

jimthompson5910 (jim_thompson5910):

then yeah, 4^2 = 16 3+1 = 4 16/4 = 4

OpenStudy (anonymous):

so its 4? im so confused i thought 4 times 2 was 8. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

4^2 means 4*4 = 16

jimthompson5910 (jim_thompson5910):

4^2 = 4 squared

OpenStudy (anonymous):

oh okay thanks, i understand now

jimthompson5910 (jim_thompson5910):

no problem